OpenStudy (anonymous):

y+x=-10 type two ordered pairs and graph the y intercept

OpenStudy (anonymous):

that is a downward straight line with angular coefficient -1 and intercept y at -10. y=-x -10

OpenStudy (anonymous):

so how would you get two ordered pairs out of that?

OpenStudy (anonymous):



OpenStudy (anonymous):

intercept on (0 , -10)

OpenStudy (anonymous):

yes that is the solution but I need something like ( , ) , ( ,) to plot tow points and get the solution.

OpenStudy (anonymous):

Getting two ordered pairs for an equation should be done as follows: 1) Pick any one value of y you like. Plug into the equation and find the corresponding value of x. Your first ordered pair will be (x you calculated, y you selected). 2) Pick any second value of y you like. Plug into the equation and find the corresponding value of x. Your second ordered pair will be (x you calculated, y you selected).

OpenStudy (anonymous):

o ok... make y=0 and solve to x 0=-x-10 x=-10 points: (-10, 0) and (0 , -10)

OpenStudy (anonymous):

put x = 0 u will get what y = -10 put y = 0 u will get x = -10 thus u having two points (0, - 10) and (-10,0)

OpenStudy (anonymous):

Thanks a million!

OpenStudy (anonymous):

Ordered pairs on a line DON'T have to be the intercepts. So, what I suggest above is a more general solution to find any two ordered pairs on the line.